# Reward Model Ensembles Mitigate but do not Eliminate Reward Hacking

### v2

@@ Abstract
Reward models play a key role in aligning language model applications towards human preferences. However, this setup creates an incentive for the language model to exploit errors in the reward model to achieve high estimated reward, a phenomenon often termed reward hacking. A natural mitigation is to train an ensemble of reward models, aggregating over model outputs to obtain a more robust reward estimate. We explore the application of reward ensembles to alignment at both training time (through reinforcement learning) and inference time (through reranking). First, we show that reward models are underspecified: reward models that perform similarly in-distribution can yield very different rewards when used in alignment, due to distribution shift. Second, underspecification results in overoptimization, where alignment to one reward model does not improve reward as measured by another reward model trained on the same data. Third, overoptimization is mitigated by the use of reward ensembles, and ensembles that vary by their pretraining seeds lead to better generalization than ensembles that differ only by their fine-tuning seeds, with both outperforming individual reward models. However, even pretrain reward ensembles do not eliminate reward hacking: we show several qualitative reward hacking phenomena that are not mitigated by ensembling because all reward models in the ensemble exhibit similar error patterns.
@@ 1 Introduction
To align machine learning systems with human preferences, it is common to use reward models that are finetuned on preference annotations to score potential outputs by how likely they are to be preferred by human raters (Christiano et al., 2017; Stiennon et al., 2020; Bai et al., 2022; Roit et al., 2023). There are many ways to use reward models to align policy models: they can act as training signals in reinforcement learning (Christiano et al., 2017; Stiennon et al., 2020), they can select examples for further imitation learning (Gulcehre et al., 2023; Liu et al., 2023; Dong et al., 2023; Touvron et al., 2023), or they can be applied at inference time to steer the output distribution toward higher expected reward (e.g., Yang & Klein, 2021; Gao et al., 2023). Such procedures create a semi-adversarial dynamic in which the language model is encouraged to produce outputs that obtain high reward by exploiting errors in the reward model. Furthermore, while the reward model is trained on a fixed set of human preference data, the process of alignment shifts the distribution of its inputs, increasing the likelihood of such errors. This phenomenon where the policy language model exploits reward model errors is often termed reward hacking (Amodei et al., 2016), reward gaming (Skalse et al., 2022; Pang et al., 2023), or reward over-optimization (Gao et al., 2023).
@@
Reward hacking has been investigated from several perspectives in prior work (e.g., Krakovna et al., 2020; Skalse et al., 2022; Pan et al., 2022). Bai et al. (2022) used reinforcement learning with human feedback (RLHF) and trained two reward models on non-overlapping splits of preference data, using one to drive alignment, and the other to measure the quality of the outputs. They find that RLHF increases performance according to both the driver and measurement models, but that a performance gap emerges as the policy is allowed to diverge from the initial distribution. However, both reward models were built on base models trained on the same pretraining data, which, as we will show, limits their diversity (as hypothesized by Gleave & Irving (2022)) and thus may understate the effect of reward hacking. Other work has simulated the relationship between a “true” reward and a learned proxy, showing that it is possible to over-optimize the proxy to such an extent that the true reward starts to decrease (Gao et al., 2023; Coste et al., 2023). This has been replicated in more realistic settings by examining (and creating) spurious correlations in reward model training data (Pang et al., 2023).
@@
In this work, we first analyze reward model distribution shift from the perspective of underspecification (D’Amour et al., 2022), which occurs when a machine learning pipeline yields reliable performance on held-out data from the training distribution, but variable performance on out-of-distribution data. When applied to learning reward models from human preference data, we show that reward models that agree in-distribution often disagree when transferred out-of-distribution. Furthermore, such disagreements are more pronounced when the reward models are built on different pretrainings, even when that difference is induced merely by varying the pretraining random seed. These disagreements become increasingly severe when evaluated on outputs of a policy model that has been aligned to a specific reward model. This occurs both when using reward models in RLHF, as well as when using an inference-time alignment procedure, best-of-$n$ reranking, where $n$ samples are drawn from the policy and then reranked with a reward model.
@@
Figure 1: Left: reward model ensembles can attenuate errors made by individual reward models, in this case the positive $r_{1}$ for this off-topic response from the policy model $\pi(y\mid x)$, which gets a low true reward ($r^{*}$). Right: insufficiently diverse reward models unanimously rate this overly-verbose and non-responsive reply from $\pi(y\mid x)$ as positive, but it too gets a low true reward. Both examples are real outputs and rewards (here, represented as normalized Z-scores) from best-of-$n$ reranking on a dataset of dialogue turns rated for helpfulness (Bai et al., 2022); see the paper for details.
@@
Motivated by these findings, we systematically investigate reward model ensembles as a possible remedy for reward hacking. Assuming different models err in different ways, ensembling can leverage reward uncertainty across the ensemble during alignment (see Figure 1, Left). We explore several techniques for aggregating scores across the ensemble, e.g., taking the median score as a robust estimate of the true reward of the policy. We also consider two types of ensembles: pretrain ensembles, where different members of the ensemble differ in the random seed used during the pretraining phase, and finetune ensembles, where members differ only in the random seed used during finetuning. These ensembles are then evaluated across several types of policies and preference annotations: dialogue preferences for a helpful assistant (Bai et al., 2022), summarization quality (Stiennon et al., 2020), and whether a single-document summary is grounded in its source text (Roit et al., 2023).
We find that pretrain ensembles substantially outperform finetune ensembles. Moreover, they consistently outperform single reward models, unlike finetune ensembles, which in many cases are comparable to single reward models. However, our analysis also reveals that policies trained with ensembles are still susceptible to reward hacking: different reward models sometimes share similar error patterns, which in turn propagate to the ensemble (see Figure 1, Right). This is exploited and amplified by the policy, leading, for example, to outputs that are too short when tuning for factuality, too verbose when tuning for summarization quality, or responses that follow a particular format that is often unsuitable, when training a helpful assistant. Thus, it is possible that methods that, unlike ensembles, are aware of the distance of outputs from the reward data distribution (Liu et al., 2020) could provide more reliable estimates of uncertainty.
In concurrent work, Coste et al. (2023) argue that reward model ensembles effectively mitigate reward hacking. Our work shares a similar research question, but differs in several ways, leading to more nuanced conclusions. First, we investigate the difference between pretrain and finetune ensembles, finding that pretrain ensembles are considerably more effective. Second, we use human-annotated preference data rather than synthetically-generated labels, which provides a more realistic experimental setup. Third, we perform analysis that demonstrates the limitations of reward ensembles, showing reward ensembles are still susceptible to reward hacking. Last, our experimental setup covers a wider range of tasks, larger reward models, and more extensive policy optimization.
@@ 2 Preliminaries
Reward models have become the primary tool for aligning LMs towards user-facing applications. We now briefly review how reward models are trained (§2.1) and how they are used for alignment (§2.2). We then describe the experimental setup that we will use for the remainder of the paper (§2.3).
@@ 2.1 Reward Model Training
We focus on the the typical setup where reward models are trained from preference data, $(x,y^{+},y^{-})\in D$, where $y^{+}$ is annotated to be preferred over $y^{-}$ for prompt $x$. Under the Bradley-Terry model (Bradley & Terry, 1952), the probability that response $y_{2}$ is preferred over $y_{1}$ given a reward function $r$ and a prompt $x$ is $p(y_{1}\prec y_{2}\mid x)=\sigma(r(x,y_{2})-r(x,y_{1}))$, where $\sigma(\cdot)$ is the sigmoid function. Then, we can use preference data to train a reward model by maximizing
@@
The Bradley-Terry model is underdetermined: for any reward model $r^{*}$, we can define an equivalent reward model, $r^{\prime}(x,y)=r^{*}(x,y)+C(x)$ where $C(x)$ is a prompt-dependent constant, obtaining the same objective value as $r^{*}$, i.e., $\mathcal{J}(r^{*})=\mathcal{J}(r^{\prime})$. This is problematic for ensembling: if different reward models choose different values for $C(x)$, then order statistics like median and minimum are meaningless. We therefore modify the objective function by adding a regularization term to encourage the sum of reward values per preference pair to stay close to zero, i.e.,
@@
Note that reward models can also be trained from “pointwise” data, such as toxicity or factuality annotations on individual examples (Yang & Klein, 2021; Roit et al., 2023). Such reward models are not underdetermined and so can be aggregated without adjustment.
@@ 2.2 Aligning Language Models using Reward Models
Best-of-$n$ reranking (BoN) is an inference-time alignment strategy, where given a prompt $x$, we sample $n$ generations $y_{1},\ldots,y_{n}$ from a policy language model $\pi(y\mid x)$ and return the generation that has the highest reward according to a reward model $r$, i.e., $y*=\arg\max_{y_{k}\in\{y_{1},\ldots,y_{n}\}}r(x,y_{k})$. The Kullback–Leibler (KL) divergence of BoN from the initial policy is upper bounded by $\log n-\frac{n-1}{n}$. BoN tends to outperform more elaborate alignment techniques like RLHF in the low-KL regime (Gao et al., 2023), albeit with the cost of generating multiple samples at inference time.
@@
where $r$ is a reward model, $\rho$ is a distribution over prompts, and $\lambda$ is a hyper-parameter. Typically, this objective is optimized using PPO (Schulman et al., 2017), which we also use in this work.
@@ 2.3 Experimental Setup
Task Prompt Output
TL;DR
I’ll try to keep this short! **Background** * I’ve always been an on again/off again (very casual!) jogger, typically doing 3 - 5 k * My knees have always been finicky, and I went to a physio who thought I had “runner’s knee” […] I had my baby a year ago, so all in all I haven’t run for about 1.5 years […] However, when I run even the tinyiest bit, or even go on a long walk or a hike, my pelvis gets very sore and tight […]
I think pregnancy messed with my body, now I can’t even run even the smallest amount without pain in my pelvis and knees. I’m fairly certain the problem isn’t just that I’m completely out of shape.
Helpfulness
Human: Assistant, can you help me find books that have really popular antagonists? Assistant: Sure! I know many books that feature good antagonists, such as those in the Lord of the Rings books. Human: Who specifically from Lord of the Rings? Assistant:
Well, Sauron, for example. If you want, I can explain who he was.
XSum/NLI
The ex-Reading defender denied fraudulent trading charges relating to the Sodje Sports Foundation - a charity to raise money for Nigerian sport. Mr Sodje, 37, is jointly charged with elder brothers Efe, 44, Bright, 50 and Stephen, 42. Appearing at the Old Bailey earlier, all four denied the offence. The charge relates to offences which allegedly took place between 2008 and 2014. Sam, from Kent, Efe and Bright, of Greater Manchester, and Stephen, from Bexley, are due to stand trial in July. They were all released on bail.
Former Premier League footballer Sam Sodje has appeared in court alongside three brothers accused of charity fraud.
We will examine the performance of reward models (both single models and ensembles) across three tasks. An example from each task is provided in Table 1.
  • [leftmargin=*]
  • tl;dr: A summarization benchmark where authors summarize their own reddit posts (Völske et al., 2017). We use the preference data created by Stiennon et al. (2020). This benchmark has been commonly used to evaluate finetuning of policy LMs (Rafailov et al., 2023; Zhao et al., 2023).
  • helpfulness: A helpful assistant benchmark (Bai et al., 2022), where given a partial conversation between a human and a digital assistant the goal is to complete the next turn of the assistant. This benchmark has also been commonly used for evaluating finetuned policy LMs (Bai et al., 2022; Rafailov et al., 2023). We use the base dataset (44K examples), where responses are generated from a 52B context-distilled LM, and split the training set into two: half for training the reward model, and half for training the policy model.
  • xsum/nli: We adopt the setup of factually-consistent summarization (Roit et al., 2023), where a model trained on XSum (Narayan et al., 2018) is finetuned to generate summaries that are consistent with the source document according to a Natural Language Inference (NLI) reward model.
@@
To examine the effect of pretraining on reward models, we pretrain five T5 models from scratch with the base (220M parameters), large (770M), and XL (3B) architectures, using the standard denoising objective over the C4 corpus (Raffel et al., 2020). The pretrained checkpoints differ only in their random seed, which controls parameter initialization and the sample from the pretraining data. The same pretrained models are used for finetuning across all tasks.
We finetune each pretrained model five times using different random seeds across all three benchmarks. In tl;dr and helpfulness we use the aforementioned preference data. For xsum/nli, we finetune NLI models on the ANLI dataset (Nie et al., 2020). Overall we obtain 25 reward models per task (5 pretrain $\times$ 5 finetune). This makes it possible to evaluate the effect of pretraining and finetuning on underspecfication (§3) by constructing ensembles that differ in either pretrain or finetune seed (§4).
Alignment strategy
We use the publicly available T5-large model (Raffel et al., 2020) as a policy for the two summarization tasks. For helpfulness, the task requires substantial background knowledge, and thus we use the instruction-tuned PALM-2-XXS model (Anil et al., 2023). Prior to alignment, we create a finetuned policy $\pi_{\text{sft}}$ by finetuning on supervised data in the standard manner. We finetune on annotated summaries from tl;dr and xsum/nli for the corresponding tasks, and on the preferred responses, $(x,y^{+})$, from the preference data in helpfulness.
In BoN reranking, we rerank sampled sets of size $n\in\{2^{1},2^{2},\ldots,2^{5}\}$ for helpfulness and $\{2^{1},\ldots,2^{6}\}$ for tl;dr. Larger sets lead to higher reward at a cost of more expensive inference and larger deviation from $\pi_{\text{sft}}$. In RLHF, we obtain a trade-off between the KL from $\pi_{\text{sft}}$ and the expected reward by training multiple times, varying the value of $\lambda$. Low values of $\lambda$ correspond to high KL and high reward, while high values of $\lambda$ entail low KL and low reward. For each value of $\lambda$ we train roughly to convergence using a predetermined fixed number of steps (all hyperparameter values, including $\lambda$ and the number of steps, are in Appendix C). Coste et al. (2023) trade-off KL and reward by tracking their values during training; however, for any particular value of KL the reward might still be underoptimized during training (i.e., there can exist a different policy $\pi(y\mid x)$ with better reward, but the same $\mathrm{KL}(\pi(y\mid x)\|\pi_{\mathrm{sft}}(y\mid x))$, which can be found with longer training).
@@
Table 2: Mean in-distribution accuracy of 25 trained reward models on validation data for tl;dr, helpfulness, and xsum/nli. Standard deviation is also reported, and observed to be small in-distribution. The single T5-XXL reward model is used for evaluation purposes only.
@@
We use two metrics to quantify generalization of reward models—reward by a larger model and win rate. Similar to past work (Gao et al., 2023; Coste et al., 2023), we use a larger reward model to evaluate the generalization of models trained with a smaller reward model. We train a T5-XXL reward model by taking the publicly available T5-XXL (Raffel et al., 2020) and finetuning it as described above. Table 2 details the performance of reward models of different sizes on the three tasks, and it can be seen that T5-XXL outperforms the best T5-XL model. We report both average reward of the T5-XXL evaluator as well as win rate, which is the fraction of prompts for which the response sampled from the aligned policy $\pi$ has higher reward compared to $\pi_{\text{sft}}$.
The errors of the T5-XXL autoeval model might correlate with errors of the smaller T5 models because they are trained on the same preference data. For this reason, we also evaluate win rate according to a prompted PaLM-2-Large model, which was not exposed to the reward training data but was instruction-tuned on FLAN (Wei et al., 2022). Given a prompt $x$, we sample a response $y_{\text{sft}}$ from $\pi_{\text{sft}}$ and $y_{\text{rlhf}}$ from $\pi$. We then ask PaLM-2 which response is better, using a hand-engineered prompt proposed by Rafailov et al. (2023). To avoid position bias we run PaLM-2 on the two possible orderings $(y_{\text{sft}},y_{\text{rlhf}})$ and $(y_{\text{sft}},y_{\text{rlhf}})$, sample $K=8$ outputs for each order and determine the winner on this prompt through majority voting. This style of evaluation has become common recently (Dubois et al., 2023; Singhal et al., 2023) and was shown to correlate well with human judgements (Rafailov et al., 2023).

### v3

# Helping or Herding? [Uncaptioned image] Reward Model Ensembles Mitigate but do not Eliminate Reward Hacking

@@ Abstract
Reward models play a key role in aligning language model applications towards human preferences. However, this setup creates an incentive for the language model to exploit errors in the reward model to achieve high estimated reward, a phenomenon often termed reward hacking. A natural mitigation is to train an ensemble of reward models, aggregating over model outputs to obtain a more robust reward estimate. We explore the application of reward ensembles to alignment at both training time (through reinforcement learning) and inference time (through reranking). First, we show that reward models are underspecified: reward models that perform similarly in-distribution can yield very different rewards when used in alignment, due to distribution shift. Second, underspecification results in overoptimization, where alignment to one reward model does not improve reward as measured by another reward model trained on the same data. Third, overoptimization is mitigated by the use of reward ensembles, and ensembles that vary by their pretraining seeds lead to better generalization than ensembles that differ only by their fine-tuning seeds, with both outperforming individual reward models. [^1] These pretrains are available at https://github.com/google-deepmind/reward-ensembles. However, even pretrain reward ensembles do not eliminate reward hacking: we show several qualitative reward hacking phenomena that are not mitigated by ensembling because all reward models in the ensemble exhibit similar error patterns.
@@ 1 Introduction
To align machine learning systems with human preferences, it is common to use reward models (RMs), which are finetuned on preference annotations to score potential outputs by how likely they are to be preferred by human raters (Christiano et al., 2017; Stiennon et al., 2020). There are many ways to use RMs to align policy models: they can act as training signals in reinforcement learning (Christiano et al., 2017; Stiennon et al., 2020), they can select examples for further imitation learning (Gulcehre et al., 2023; Liu et al., 2023; Dong et al., 2023; Touvron et al., 2023), or they can be applied at inference time to steer the output distribution toward higher expected reward (e.g., Yang & Klein, 2021; Gao et al., 2023). Such procedures create a semi-adversarial dynamic, in which the language model can achieve high reward by exploiting errors in the RM. Furthermore, while the RM is trained on a fixed set of human preference data, the process of alignment shifts the distribution of its inputs, increasing the likelihood of such errors. This phenomenon where the policy language model exploits reward model errors is often termed reward hacking (Amodei et al., 2016), reward gaming (Skalse et al., 2022; Pang et al., 2023), or reward over-optimization (Gao et al., 2023).
@@
Reward hacking has been investigated from several perspectives in prior work (e.g., Krakovna et al., 2020; Skalse et al., 2022; Pan et al., 2022). Bai et al. (2022) used reinforcement learning with human feedback (RLHF) and trained two RMs on non-overlapping splits of preference data, using one to drive alignment, and the other to measure the quality of the outputs. They find that RLHF increases performance according to both the driver and measurement models, but that a performance gap emerges as the policy is allowed to diverge from the initial distribution. However, both RMs were built on base models trained on the same pretraining data, which, as we will show, limits their diversity (as hypothesized by Gleave & Irving (2022)) and thus may understate the effect of reward hacking. Other work has simulated the relationship between a “true” reward and a learned proxy, showing that it is possible to over-optimize the proxy to such an extent that the true reward starts to decrease (Gao et al., 2023; Coste et al., 2023), e.g. by exploiting spurious correlations in reward model training data (Pang et al., 2023).
@@
In this work, we first analyze RM distribution shift from the perspective of underspecification (D’Amour et al., 2022), which occurs when a machine learning pipeline yields reliable performance on held-out data from the training distribution, but variable performance on out-of-distribution data. When applied to learning RMs from human preference data, we show that RMs that agree in-distribution often disagree after alignment-induced policy distribution shifts. Furthermore, such disagreements are more pronounced when the RMs are built on different pretrainings, even when that difference is induced merely by varying the pretraining random seed. These disagreements become increasingly severe when evaluated on outputs of a policy model that has been aligned to a specific RM. This occurs both when using RMs in RLHF, as well as when using an inference-time alignment procedure, best-of-$n$ reranking, where $n$ samples are drawn from the policy and then reranked with a RM.
@@
Figure 1: Left: RM ensembles can attenuate errors made by individual RMs, in this case the positive $r_{1}$ for this off-topic response from the policy model $\pi(y\mid x)$, which gets a low true reward ($r^{*}$). Right: insufficiently diverse RMs unanimously rate this overly-verbose and non-responsive reply from $\pi(y\mid x)$ as positive, but it too gets a low true reward. Both examples are real outputs and rewards (represented as normalized Z-scores) from RMs trained on the helpfulness data of Bai et al. (2022).
@@
Motivated by these findings, we systematically investigate reward model ensembles as a possible remedy for reward hacking. Assuming different models err in different ways, ensembling can leverage reward uncertainty across the ensemble during alignment (see Figure 1, Left). We explore several techniques for aggregating scores across the ensemble, e.g., taking the median score as a robust estimate of the true reward of the policy. We also consider two types of ensembles: pretrain ensembles, where different members of the ensemble differ in the random seed used during the pretraining phase, and finetune ensembles, where members differ only in the random seed used during finetuning. These ensembles are then evaluated across several types of policies and preference annotations: dialogue preferences for a helpful assistant (Bai et al., 2022), summarization quality (Stiennon et al., 2020), and whether a summary is grounded in its source text (Roit et al., 2023).
We find that ensembles are significantly more robust than individual reward models, and that pretrain ensembles are particuarly effective. However, reward ensembles are still susceptible to reward hacking when all members of the ensemble share similar error patterns, which in turn propagate to the ensemble (see Figure 1, Right). This is exploited and amplified by policy optimization: for example, summarization models produce outputs that are too short when tuned for factuality, but too verbose when tuned for summarization quality; assistant models overuse formulaic answer formats when tuned for helpfulness.
Recent and concurrent related work
Coste et al. (2023) argue that reward model ensembles effectively mitigate reward hacking. Our work shares a similar research question, but differs in several ways, leading to more nuanced conclusions: we investigate both pretrain and finetune ensembles; we use human-annotated preference data rather than synthetically-generated labels; and we analyze several real reward hacks, showing that they are not prevented by ensembles. Subsequent work has explored efficient RM ensembles through low-rank adaptation (e.g., Zhai et al., 2023) and weight averaging (Ramé et al., 2024), but exclusively for finetune ensembles. More generally, we can try to optimize the policy against the worst-case reward in some analytically-computed uncertainty set (Zhu et al., 2023; Zhang et al., 2024), but this too accounts only for uncertainty due to the finite sample of preference pairs and not in the representations made available through pretraining.
@@ 2 Preliminaries
We now briefly review how reward models are trained (§2.1) and how they are used for alignment (§2.2). We then describe our setup for studying reward hacking (§2.3).
@@ 2.1 Reward Model Training
RMs are typically trained from preference data, $(x,y^{+},y^{-})\in D$, where $y^{+}$ is preferred over $y^{-}$ for prompt $x$. Under the Bradley-Terry model (Bradley & Terry, 1952), the probability that response $y_{2}$ is preferred over $y_{1}$ given a reward function $r$ and a prompt $x$ is $p(y_{1}\prec y_{2}\mid x)=\sigma(r(x,y_{2})-r(x,y_{1}))$, where $\sigma(\cdot)$ is the sigmoid function. Given a dataset of preferences, the maximum-likelihood objective is,
@@
The Bradley-Terry model is underdetermined: for any RM $r^{*}$, there is a set of equivalent models, $r^{\prime}(x,y)=r^{*}(x,y)+C(x)$ where $C(x)$ is a prompt-dependent constant, such that $\mathcal{J}(r^{*})=\mathcal{J}(r^{\prime})$. This is problematic for ensembling: if different RMs choose different values for $C(x)$, then order statistics like median and minimum are meaningless. We therefore modify the objective by regularizing the sum of rewards per preference pair towards zero:
@@
Note that RMs can also be trained from “pointwise” data, such as toxicity or factuality annotations on individual examples (Yang & Klein, 2021; Roit et al., 2023). Such RMs are not underdetermined and so can be aggregated without adjustment.
@@ 2.2 Aligning Language Models using Reward Models
Best-of-$n$ reranking (BoN) is an inference-time alignment strategy, where given a prompt $x$, we sample $n$ generations $y_{1},\ldots,y_{n}$ from a policy language model $\pi(y\mid x)$ and return the generation with highest reward: $y*=\arg\max_{y_{k}\in\{y_{1},\ldots,y_{n}\}}r(x,y_{k})$. The Kullback–Leibler (KL) divergence of BoN from the initial policy is upper bounded by $\log n-\frac{n-1}{n}$ (Beirami et al., 2024). BoN tends to outperform more elaborate alignment techniques like RLHF in the low-KL regime (Gao et al., 2023), albeit with the cost of generating multiple samples at inference time.
@@
where $r$ is an RM, $\rho$ is a distribution over prompts, and $\lambda$ is a hyperparameter. Typically, this objective is optimized using PPO (Schulman et al., 2017), which we also use in this work.
The KL penalty in eq. 3 can limit reward hacking by keeping the policy close to $\pi_{\text{sft}}.$ However, KL regularization does not directly address reward model errors, and in particular does not address RM distribution shift when the preference annotations are not sampled from $\pi_{\text{sft}}$ (as is generally the case). Furthermore, we want to diverge from the reference policy when we can be confident of improving the expected reward. As we will show empirically, KL regularization is complementary to the development of more robust reward models, which yields pareto improvements in the reward-KL tradeoff.
@@ 2.3 Experimental Setup
We consider three tasks. Example instances are provided in Table 12.
  • tl;dr: A summarization benchmark where authors summarize their own reddit posts (Völske et al., 2017; Stiennon et al., 2020), frequently used in research on RLHF and related methods (Rafailov et al., 2023; Zhao et al., 2023).
  • helpfulness: A popular dialogue benchmark (Bai et al., 2022), where given a partial conversation between a human and a digital assistant the goal is to complete the next assistant turn (Bai et al., 2022; Rafailov et al., 2023). [^2] We use the base dataset (44K examples), where responses are generated from a 52B context-distilled LM, and split the training set into two: half for training the RM, and half for the policy.
  • xsum/nli: A summarization benchamark where a policy model trained on XSum (Narayan et al., 2018) is finetuned to generate summaries that are factually consistent with the source document (Roit et al., 2023).
@@
To examine the effect of pretraining, we pretrain five T5 models from scratch at the base (220M parameters), large (770M), and XL (3B) scales, using the standard denoising objective over the C4 corpus (Raffel et al., 2020). The pretrained checkpoints differ only in their random seed, which controls parameter initialization and the sample from the pretraining data.
For each task, we finetune each pretrained model five times using different random seeds. In tl;dr and helpfulness we use the aforementioned preference data. For xsum/nli, we finetune pointwise natural language inference (NLI) models on the ANLI dataset (Nie et al., 2020). This yields 25 RMs per task and scale (5 pretrain $\times$ 5 finetune), making it possible to evaluate the effect of pretraining and finetuning on underspecification (§3) by constructing ensembles that differ in either pretrain or finetune seed (§4).
@@
Table 1: Mean in-distribution accuracy of 25 RMs on validation data for tl;dr, helpfulness, and xsum/nli. Standard deviation, indicated with $\pm$, is small in-distribution. The single T5-XXL RM is used for evaluation purposes only.
Alignment strategy
We use the publicly available T5-large (Raffel et al., 2020) as a policy for the summarization tasks. For helpfulness, which requires substantial background knowledge, we use the instruction-tuned PALM-2-XXS model (Anil et al., 2023). Before alignment, we create a finetuned policy $\pi_{\text{sft}}$ through supervised finetuning: We finetune on annotated summaries from tl;dr and xsum/nli for the corresponding tasks, and on the preferred responses, $(x,y^{+})$, from the preference data in helpfulness.
In BoN reranking, we rerank sampled output sets of size $n\in\{2^{1},2^{2},\ldots,2^{5}\}$ for helpfulness and $\{2^{1},\ldots,2^{6}\}$ for tl;dr. Larger sets lead to higher reward at a cost of more expensive inference and larger deviation from $\pi_{\text{sft}}$. In RLHF, we obtain a trade-off between the KL from $\pi_{\text{sft}}$ and the expected reward by training multiple times, varying the value of $\lambda$. Low values of $\lambda$ correspond to high KL and high reward, while high values of $\lambda$ entail low KL and low reward. For each value of $\lambda$ we train roughly to convergence using a predetermined fixed number of steps (all hyperparameter values, including $\lambda$ and the number of steps, are in Appendix C). Coste et al. (2023) trade-off KL and reward by tracking their values during training; however, for any particular value of KL the reward might still be underoptimized during training (i.e., there can exist a different policy $\pi(y\mid x)$ with better reward, but the same $\mathrm{KL}(\pi(y\mid x)\|\pi_{\mathrm{sft}}(y\mid x))$, which can be found with longer training).
@@
The aligned policies are autoevaluated by two metrics: reward and win rate. Similar to past work (Gao et al., 2023; Coste et al., 2023), we use a larger RM to evaluate the generalization of models trained with a smaller RM. We train a T5-xxl RM by taking the publicly available T5-xxl (Raffel et al., 2020) and finetuning it as described above. As shown in Table 1, T5-xxl outperforms the best T5-xl model on each task. We report both average reward from the T5-xxl evaluator as well as win rate, which is the fraction of prompts for which the response sampled from the aligned policy $\pi$ has higher reward compared to $\pi_{\text{sft}}$.
Because the T5-xxl autoeval model is trained on the same data as the smaller T5 RMs, their errors might be correlated. For this reason, we also compute win rate according to a prompted PaLM-2-Large model, which was not exposed to the reward training data but was instruction-tuned on Flan (Wei et al., 2022). Given a prompt $x$, we sample a response $y_{\text{sft}}$ from $\pi_{\text{sft}}$ and $y_{\text{rlhf}}$ from $\pi$. We then ask PaLM-2 which response is better, using a hand-engineered prompt proposed by Rafailov et al. (2023). To avoid position bias, we run PaLM-2 on the two possible orderings $(y_{\text{sft}},y_{\text{rlhf}})$ and $(y_{\text{sft}},y_{\text{rlhf}})$, sample $K=8$ outputs for each order and determine the winner on this prompt through majority voting. This style of evaluation has become popular (Dubois et al., 2023; Singhal et al., 2023) and was shown to correlate well with human judgements (Rafailov et al., 2023).